# Ambiguous Online Learning

### v3

@@ Appendix C Mistakes and the Ambiguous Littlestone Dimension
Consider any learner $A:(X\times Y)^{*}\times X\to 2^{Y}$. Construct $xy\in(X\times Y)^{*}$ and $v\in V^{*}$ recursively as follows. We start from $v_{0}$. Suppose we constructed $xy_{(:k)}$ and $v_{(:(k+1))}$ for some $k\in\mathbb{N}$. If $v_{k}\in V_{f}$, the process ends (i.e. $|xy|=k$ and $|v|=k+1$). Otherwise, set $x_{k}:=\mathbf{x}_{v_{k}}$ and consider $\alpha_{k}:=A(xy_{(:k)},x_{k})$. If there is some $a\in P^{-1}(v_{k})$ s.t. $\mathbf{y}_{a}\notin\alpha_{k}$, set $y_{k}:=\mathbf{y}_{a}$ and $v_{k+1}:=a$. Otherwise, let $b$ be the unique element of $E_{0}\cap P^{-1}(v_{k})$. Set $y_{k}:=\mathbf{y}_{b}$ and $v_{k+1}:=b$.
@@
We conclude that $\mathcal{M}_{\mathcal{H}}^{A}(n)\geq\mathcal{M}_{h}^{A}(n)\geq\mathcal{M}_{h}^{A}(m)\geq|R_{T}(v_{m})|\geq r$. Since this holds for any AA, we get $\mathcal{M}_{\mathcal{H}}^{*}(n)\geq r$. ∎
@@
of Proposition 4.7.
Consider any $N\in\mathbb{N}$ and $r\leq\mathrm{AL}(\mathcal{H},N)$. By Proposition 3.10, there exists an ambiguous shattered $\mathcal{H}$-tree $T$ with $\operatorname{dep}(T)\leq N$, $\operatorname{rank}(T)\geq r$ and $\operatorname{ar}(T)\leq\overline{\mathrm{PD}}(\mathcal{H})+1$. Denote $q:=\operatorname{ar}(T)$.
Consider any randomized learner $R:(X\times Y)^{*}\times X\to\Delta 2^{Y}$. Construct $xy\in(X\times Y)^{*}$ and $v\in V^{*}$ recursively as follows. We start from $v_{0}$. Suppose we constructed $xy_{(:k)}$ and $v_{(:(k+1))}$ for some $k\in\mathbb{N}$. If $v_{k}\in V_{f}$, the process ends (i.e. $|xy|=k$ and $|v|=k+1$). Otherwise, set $x_{k}:=\mathbf{x}_{v_{k}}$ and consider $\theta_{k}:=R(xy_{(:k)},x_{k})$. Define
@@
Here, AA is some learner and $w^{\overline{xy}}$ is a function from $\mathcal{H}_{\text{uf}}^{\overline{xy}}$ to $\mathbb{N}$.
@@ Appendix D Trimming Shattered Trees
The idea of the proof is judiciously selecting NN edges out of the $E_{0}$-path starting at the root, and then using induction to trim the subtrees attached to these edges.
of Lemma D.1.
We proceed by induction on NN.
@@
Now, assume the claim holds for some NN. Consider an $\mathcal{H}$-tree $T=(V,v_{0},P,\mathbf{x},\mathbf{y},\mathbf{h},E_{0})$ s.t. $\operatorname{rank}(T)\geq(N+1)^{2}$. Construct the sequence $\{v_{i}\in V\}_{i\leq d}$, for some $d\in\mathbb{N}$, as follows:
@@ Appendix E Mistake Bound for WAA
The following lemma shows that the rate with which $Z$ grows along an $\mathcal{H}$-realizable trace can be bounded. Here’s the idea of the proof. If the new label $y^{*}$ is outside the predicted set $\alpha^{*}$, then a substantial fraction (measured by $\mu^{w}$) of hypotheses become falsified: if a lot of hypotheses were consistent with $y^{*}$, the learner would assign $y^{*}$ into its prediction. If $y^{*}$ is inside $\alpha^{*}$, then the fraction of hypotheses that gain a mistake cannot be too great. Because, we can choose a set $\beta\subseteq\alpha^{*}$ of “pivots” whose cardinality is at most $\underline{\mathrm{PD}}(\mathcal{H})$, and only hypotheses that exclude at least one of the pivots gain a mistake, but each pivot has to be contained in a substantial fraction of hypotheses since it was selected into $\alpha^{*}$.
@@
$Z(\overline{xy}x^{*}y^{*})\leq\max(\mu\nu,1+\underline{\mathrm{PD}}(\mathcal{H})\cdot(\mu-1)(1-\nu))Z(\overline{xy})$ (11)
@@
Now, consider the case $y^{*}\in\alpha^{*}$. Choose $\beta\subseteq\alpha^{*}$ s.t. $\alpha^{*}\subseteq\beta^{(\Lambda(\mathcal{H}))}$ and $|\beta|\leq\underline{\mathrm{PD}}(\mathcal{H})$. Then, for any $h\in\mathcal{H}$, if $\beta\subseteq h(x^{*})$ then $\alpha^{*}\subseteq h(x^{*})$. Such a hypothesis $h$ doesn’t gain a mistake. Conversely, if $\alpha^{*}\setminus h(x^{*})\neq\emptyset$ then $\beta\setminus h(x^{*})\neq\emptyset$. Denote
@@
$\leq(1-\nu)\underline{\mathrm{PD}}(\mathcal{H})Z(\overline{xy})$
@@
$\leq Z(\overline{xy})+(\mu-1)(1-\nu)\underline{\mathrm{PD}}(\mathcal{H})Z(\overline{xy})$
$=(1+\underline{\mathrm{PD}}(\mathcal{H})\cdot(\mu-1)(1-\nu))Z(\overline{xy})$
We are now ready to complete the proof of Theorem 6.1. In order to show inequality (2), we apply Lemma E.1 on each of the NN rounds, for the value of $\nu$ that corresponds to equation (1). Equation (3) follows from the inequality (2) when we choose appropriate values of $\mu$.
of Theorem 6.1.
Denote $D:=\underline{\mathrm{PD}}(\Lambda(\mathcal{H}))$. For the WAA, we have
@@ Appendix F Reduction to Finite Classes
In this section we prove Theorem $4.$
@@
$S_{n+1}:=\begin{cases}S_{n}&\text{if }A_{\text{SOA}}(h_{f_{n}}^{*}(x_{(:n)}),x_{n})=h(x_{n}),\\ S_{n}\cup\{n\}&\text{otherwise}\end{cases}$
@@
of Theorem $4.$
Consider any $N\in\mathbb{N}$, $h\in\mathcal{H}$ and $xy\in\mathcal{C}_{h}(N)$. For all $n<N$, $y_{n}\in h(x_{n})$ and in particular $h(x_{n})\neq\emptyset$. By Lemma 6.2, there exists $h^{\prime}\in\mathcal{H}^{(N)}$ s.t. for all $n<N$, $h^{\prime}(x_{(:n+1)})=h(x_{n})$. Define $\tilde{x}\in(X^{(N)})^{N}$ by $\tilde{x}_{n}:=x_{(:n+1)}$. Then, $\tilde{x}y\in\mathcal{C}_{h^{\prime}}(N)$. By Theorem 6.1,
$|\mathcal{M}^{A_{\text{WAA}}}_{h^{\prime}}(\tilde{x}y)|\leq C_{1}\cdot\underline{\mathrm{PD}}(\mathcal{H})\sqrt{N\log|\mathcal{H}^{(N)}|}$
@@
$|\mathcal{M}^{A_{\text{WAA}}}_{h^{\prime}}(\tilde{x}y)|\leq C_{1}\sqrt{C_{2}}\cdot\underline{\mathrm{PD}}(\mathcal{H})\sqrt{\mathrm{L}_{\mathfrak{P}}(\mathcal{H})N\log(|\Lambda(\mathcal{H})|N)}$
Now, consider the learner $A^{\prime}_{\text{WAA}}$ defined by $A^{\prime}_{\text{WAA}}(xy_{(:n)},x_{n}):=A_{\text{WAA}}((\tilde{x}y)_{(:n)},\tilde{x}_{n})$. We have $\mathcal{M}^{A_{\text{WAA}}}_{h^{\prime}}(\tilde{x}y)=\mathcal{M}^{A^{\prime}_{\text{WAA}}}_{h}(xy)$ and hence $A^{\prime}_{\text{WAA}}$ satisfies the required mistake bound. Indeed, for any $n<N$, the prediction of $A^{\prime}_{\text{WAA}}$ for instance $x_{n}$ is based on the prediction of $A_{\text{WAA}}$ for instance $\tilde{x}_{n}$, and $h(x_{n})=h^{\prime}(\tilde{x}_{n})$, so the corresponding mistake conditions coincide. ∎
@@ Appendix G Apple Tasting Details
First, let’s formally define the apple tasting setting. Let $X$ be the domain. Given a learner $A:(X\times\{0,1\})^{*}\times X\to\{0,1\}$, a hypothesis $h:X\to\{0,1\}$, some $n\in\mathbb{N}$ and $\in(X\times\{0,1\})^{n}, xyxy$ is said to be compatible with $h$ and AA when for all $k<n$, if $A(xy_{(:k)},x_{k})=0$ then $y_{k}=1$ and if $A(xy_{(:k)},x_{k})=1$ then $y_{k}=h(x_{k})$. Here, $y$ represents the feedbacks received by the learner (rather than the true labels). We denote by $\mathcal{C}_{h}^{A}(n)$ the set of all such xyxy.
Given $xy\in\mathcal{C}_{h}^{A}(n)$, we define the set of mistakes that AA makes on xyxy relative to $h$ by
$\mathcal{PM}_{h}^{A}(xy):=\{k<n\mid A(xy_{(:k)},x_{k})\neq h(x_{k})\}$
@@
We now come to the proof of Proposition 7.1. Informally, we’re comparing the apple tasting learner $A^{\text{ap}}$ interacting with an environment that obeys $h$ to the ambiguous learner AA interacting with an environment that obeys $h^{\text{am}}$. We argue that erroneously predicting 0 in apple tasting corresponds to an underconfidence mistake in ambiguous online learning, whereas erroneously predicting 11 in apple tasting corresponds to an overconfidence mistake in ambiguous online learning. There are two potentially substantial differences. First, in apple tasting we receive no feedback when we predict 0, whereas in ambiguous online learning we still receive a label. However, in the latter setting, the adversary can always choose the label 11, and when the learner predicts $\{0,1\}$ there is no incentive for the adversary to choose 0. Hence, effectively there is no useful feedback in the ambiguous case as well. Second, in apple tasting we receive full information about $h(x)$ when we predict 11, whereas in ambiguous online learning the adversary may choose the label 11 even if $h(x)=\{0,1\}$ (but in this case we incur no mistake). However, this only gives the adversary additional options, so the ambiguous version can only be harder.
of Proposition 7.1.
Consider any $xy\in\mathcal{C}_{h}^{A^{\text{ap}}}(N)$. For any $k<N$, there are two cases: $A^{\text{ap}}(xy_{(:k)},x_{k})=0$ and $A^{\text{ap}}(xy_{(:k)},x_{k})=1$. If $A^{\text{ap}}(xy_{(:k)},x_{k})=0$ then $y_{k}=1$, and in particular $y_{k}\in h^{\text{am}}(x_{k})$, since $1\in h^{\text{am}}(x)$ for all $x\in X$. If $A^{\text{ap}}(xy_{(:k)},x_{k})=1$ then $y_{k}=h(x_{k})$, and in particular $y_{k}\in h^{\text{am}}(x_{k})$, since $\{h(x)\}\subseteq h^{\text{am}}(x)$ for all $x\in X$. In either case $y_{k}\in h^{\text{am}}(x_{k})$. Hence, $xy\in\mathcal{C}_{h^{\text{am}}}(N)$.
Consider any $k\in\mathcal{PM}_{h}^{A^{\text{ap}}}(xy)$. Again, there are two cases: $A^{\text{ap}}(xy_{(:k)},x_{k})=0$ and $A^{\text{ap}}(xy_{(:k)},x_{k})=1$.
If $A^{\text{ap}}(xy_{(:k)},x_{k})=0$ then $h(x_{k})=1$ and $0\in A(xy_{(:k)},x_{k})$. It follows that $h^{\text{am}}(x_{k})=\{1\}$ and $0\in A(xy_{(:k)},x_{k})\setminus h^{\text{am}}(x_{k})$. In particular, $A(xy_{(:k)},x_{k})\not\subseteq h^{\text{am}}(x_{k})$ and hence $k\in\mathcal{M}_{h^{\text{am}}}^{A}(xy)$.
If $A^{\text{ap}}(xy_{(:k)},x_{k})=1$ then $h(x_{k})=0$ and $0\notin A(xy_{(:k)},x_{k})$. Hence, $y_{k}=0$, because $xy\in\mathcal{C}_{h}^{A^{\text{ap}}}(N)$. In particular, $y_{k}\notin A(xy_{(:k)},x_{k})$ and hence $k\in\mathcal{M}_{h^{\text{am}}}^{A}(xy)$.
@@
Now, let’s prove Proposition 7.2. The idea is that any frugal ambiguous shattered tree can be interpreted as an adversary strategy for the apple tasting setting. The strategy works by going down the 0 edge when the learner predicts 11 and going down the 11 edge (which has to be the $E_{0}$ edge) when the learner predicts 0. In the former case, we always get a mistake (and the edge is always relevant). In the latter case, we get a mistake if and only if the edge is relevant to the ultimate leaf.
of Proposition 7.2.
@@
Construct $xy\in(X\times\{0,1\})^{*}$ and $v\in V^{*}$ recursively as follows. We start from $v_{0}$. Suppose we constructed $xy_{(:k)}$ and $v_{(:(k+1))}$ for some $k\in\mathbb{N}$. If $v_{k}\in V_{f}$, the process ends (i.e. $|xy|=k$ and $|v|=k+1$). Otherwise, set $x_{k}:=\mathbf{x}_{v_{k}}$. Let $z_{k}:=A(xy_{(:k)},x_{k})$. Set $y_{k}:=1-z_{k}$ and $v_{k+1}:=C_{y_{k}}(v_{k})$.

### v4

@@ Appendix C Mistakes and the Ambiguous Littlestone Dimension
Consider any learner $A:(X\times Y)^{*}\times X\to 2^{Y}$. Construct $xy\in(X\times Y)^{*}$ and $v\in V^{*}$ recursively as follows. We start from $v_{0}$. Suppose we constructed $xy_{:k}$ and $v_{:(k+1)}$ for some $k\in\mathbb{N}$. If $v_{k}\in V_{f}$, the process ends (i.e. $|xy|=k$ and $|v|=k+1$). Otherwise, set $x_{k}:=\mathbf{x}_{v_{k}}$ and consider $\alpha_{k}:=A(xy_{:k},x_{k})$. If there is some $a\in P^{-1}(v_{k})$ s.t. $\mathbf{y}_{a}\notin\alpha_{k}$, set $y_{k}:=\mathbf{y}_{a}$ and $v_{k+1}:=a$. Otherwise, let $b$ be the unique element of $E_{0}\cap P^{-1}(v_{k})$. Set $y_{k}:=\mathbf{y}_{b}$ and $v_{k+1}:=b$.
@@
We conclude that $\mathcal{M}_{\mathcal{H}}^{A}(n)\geq\mathcal{M}_{h}^{A}(n)\geq\mathcal{M}_{h}^{A}(m)\geq|R_{T}(v_{m})|\geq r$. Since this holds for any $A$, we get $\mathcal{M}_{\mathcal{H}}^{*}(n)\geq r$. ∎
@@
Proof of Proposition 4.7.
Consider any $N\in\mathbb{N}$ and $r\leq\mathrm{AL}(\mathcal{H},N)$. By Proposition 3.10, there exists an ambiguous shattered $\mathcal{H}$-tree $T$ with $\operatorname{dep}(T)\leq N$, $\operatorname{rank}(T)\geq r$ and $\operatorname{ar}(T)\leq\mathrm{VC}(\Lambda(\mathcal{H}))+1$. Denote $q:=\operatorname{ar}(T)$.
Consider any randomized learner $R:(X\times Y)^{*}\times X\to\Delta 2^{Y}$. Construct $xy\in(X\times Y)^{*}$ and $v\in V^{*}$ recursively as follows. We start from $v_{0}$. Suppose we constructed $xy_{:k}$ and $v_{:(k+1)}$ for some $k\in\mathbb{N}$. If $v_{k}\in V_{f}$, the process ends (i.e. $|xy|=k$ and $|v|=k+1$). Otherwise, set $x_{k}:=\mathbf{x}_{v_{k}}$ and consider $\theta_{k}:=R(xy_{:k},x_{k})$. Define
@@
Here, $A$ is some learner and $w^{\overline{xy}}$ is a function from $\mathcal{H}_{\text{uf}}^{\overline{xy}}$ to $\mathbb{N}$.
@@ Appendix D Trimming Shattered Trees
The idea of the proof is judiciously selecting $N$ edges out of the $E_{0}$-path starting at the root, and then using induction to trim the subtrees attached to these edges.
Proof of Lemma D.1.
We proceed by induction on $N$.
@@
Now, assume the claim holds for some $N$. Consider an $\mathcal{H}$-tree $T=(V,v_{0},P,\mathbf{x},\mathbf{y},\mathbf{h},E_{0})$ s.t. $\operatorname{rank}(T)\geq(N+1)^{2}$. Construct the sequence $\{v_{i}\in V\}_{i\leq d}$, for some $d\in\mathbb{N}$, as follows:
@@ Appendix E Mistake Bound for WAA
The following lemma shows that the rate with which $Z$ grows along an $\mathcal{H}$-realizable trace can be bounded. Here’s the idea of the proof. If the new label $y^{*}$ is outside the predicted set $\alpha^{*}$, then a substantial fraction (measured by $\mu^{w}$) of hypotheses become falsified: if a lot of hypotheses were consistent with $y^{*}$, the learner would assign $y^{*}$ into its prediction. If $y^{*}$ is inside $\alpha^{*}$, then the fraction of hypotheses that gain a mistake cannot be too great. Because, we can choose a set $\beta\subseteq\alpha^{*}$ of “pivots” whose cardinality is at most $\mathrm{PD}(\mathcal{H})$, and only hypotheses that exclude at least one of the pivots gain a mistake, but each pivot has to be contained in a substantial fraction of hypotheses since it was selected into $\alpha^{*}$.
@@
$Z(\overline{xy}x^{*}y^{*})\leq\max(\mu\nu,1+\mathrm{PD}(\mathcal{H})\cdot(\mu-1)(1-\nu))Z(\overline{xy})$ (11)
@@
Now, consider the case $y^{*}\in\alpha^{*}$. Choose $\beta\subseteq\alpha^{*}$ s.t. $\alpha^{*}\subseteq\beta^{(\Lambda(\mathcal{H}))}$ and $|\beta|\leq\mathrm{PD}(\mathcal{H})$. Then, for any $h\in\mathcal{H}$, if $\beta\subseteq h(x^{*})$ then $\alpha^{*}\subseteq h(x^{*})$. Such a hypothesis $h$ doesn’t gain a mistake. Conversely, if $\alpha^{*}\setminus h(x^{*})\neq\emptyset$ then $\beta\setminus h(x^{*})\neq\emptyset$. Denote
@@
$\leq(1-\nu)\mathrm{PD}(\mathcal{H})Z(\overline{xy})$
@@
$\leq Z(\overline{xy})+(\mu-1)(1-\nu)\mathrm{PD}(\mathcal{H})Z(\overline{xy})$
$=(1+\mathrm{PD}(\mathcal{H})\cdot(\mu-1)(1-\nu))Z(\overline{xy})$
We are now ready to complete the proof of Theorem 6.1. In order to show inequality (2), we apply Lemma E.1 on each of the $N$ rounds, for the value of $\nu$ that corresponds to equation (1). Equation (3) follows from the inequality (2) when we choose appropriate values of $\mu$.
Proof of Theorem 6.1.
Denote $D:=\mathrm{PD}(\Lambda(\mathcal{H}))$. For the WAA, we have
@@ Appendix F Reduction to Finite Classes
In this section we prove Theorem 4.4.
@@
$S_{n+1}:=\begin{cases}S_{n}&\text{if }A_{\text{SOA}}(h_{f_{n}}^{*}(x_{:n}),x_{n})=h(x_{n}),\\ S_{n}\cup\{n\}&\text{otherwise}\end{cases}$
@@
Proof of Theorem 4.4.
Consider any $N\in\mathbb{N}$, $h\in\mathcal{H}$ and $xy\in\mathcal{C}_{h}(N)$. For all $n<N$, $y_{n}\in h(x_{n})$ and in particular $h(x_{n})\neq\emptyset$. By Lemma 6.2, there exists $h^{\prime}\in\mathcal{H}^{(N)}$ s.t. for all $n<N$, $h^{\prime}(x_{:n+1})=h(x_{n})$. Define $\tilde{x}\in(X^{(N)})^{N}$ by $\tilde{x}_{n}:=x_{:n+1}$. Then, $\tilde{x}y\in\mathcal{C}_{h^{\prime}}(N)$. By Theorem 6.1,
$|\mathcal{M}^{A_{\text{WAA}}}_{h^{\prime}}(\tilde{x}y)|\leq C_{1}\cdot\mathrm{PD}(\mathcal{H})\sqrt{N\log|\mathcal{H}^{(N)}|}$
@@
$|\mathcal{M}^{A_{\text{WAA}}}_{h^{\prime}}(\tilde{x}y)|\leq C_{1}\sqrt{C_{2}}\cdot\mathrm{PD}(\mathcal{H})\sqrt{\mathrm{L}_{\mathfrak{P}}(\mathcal{H})N\log(|\Lambda(\mathcal{H})|N)}$
Now, consider the learner $A^{\prime}_{\text{WAA}}$ defined by $A^{\prime}_{\text{WAA}}(xy_{:n},x_{n}):=A_{\text{WAA}}(\tilde{x}y_{:n},\tilde{x}_{n})$. We have $\mathcal{M}^{A_{\text{WAA}}}_{h^{\prime}}(\tilde{x}y)=\mathcal{M}^{A^{\prime}_{\text{WAA}}}_{h}(xy)$ and hence $A^{\prime}_{\text{WAA}}$ satisfies the required mistake bound. Indeed, for any $n<N$, the prediction of $A^{\prime}_{\text{WAA}}$ for instance $x_{n}$ is based on the prediction of $A_{\text{WAA}}$ for instance $\tilde{x}_{n}$, and $h(x_{n})=h^{\prime}(\tilde{x}_{n})$, so the corresponding mistake conditions coincide. ∎
@@ Appendix G Apple Tasting Details
First, let’s formally define the apple tasting setting. Let $X$ be the domain. Given a learner $A:(X\times\{0,1\})^{*}\times X\to\{0,1\}$, a hypothesis $h:X\to\{0,1\}$, some $n\in\mathbb{N}$ and $\in(X\times\{0,1\})^{n}, xyxy$ is said to be compatible with $h$ and $A$ when for all $k<n$, if $A(xy_{:k},x_{k})=0$ then $y_{k}=1$ and if $A(xy_{:k},x_{k})=1$ then $y_{k}=h(x_{k})$. Here, $y$ represents the feedbacks received by the learner (rather than the true labels). We denote by $\mathcal{C}_{h}^{A}(n)$ the set of all such $xy$.
Given $xy\in\mathcal{C}_{h}^{A}(n)$, we define the set of mistakes that $A$ makes on $xy$ relative to $h$ by
$\mathcal{PM}_{h}^{A}(xy):=\{k<n\mid A(xy_{:k},x_{k})\neq h(x_{k})\}$
@@
We now come to the proof of Proposition 7.1. Informally, we’re comparing the apple tasting learner $A^{\text{ap}}$ interacting with an environment that obeys $h$ to the ambiguous learner $A$ interacting with an environment that obeys $h^{\text{am}}$. We argue that erroneously predicting 0 in apple tasting corresponds to an underconfidence mistake in ambiguous online learning, whereas erroneously predicting $1$ in apple tasting corresponds to an overconfidence mistake in ambiguous online learning. There are two potentially substantial differences. First, in apple tasting we receive no feedback when we predict 0, whereas in ambiguous online learning we still receive a label. However, in the latter setting, the adversary can always choose the label $1$, and when the learner predicts $\{0,1\}$ there is no incentive for the adversary to choose 0. Hence, effectively there is no useful feedback in the ambiguous case as well. Second, in apple tasting we receive full information about $h(x)$ when we predict $1$, whereas in ambiguous online learning the adversary may choose the label $1$ even if $h(x)=\{0,1\}$ (but in this case we incur no mistake). However, this only gives the adversary additional options, so the ambiguous version can only be harder.
Proof of Proposition 7.1.
Consider any $xy\in\mathcal{C}_{h}^{A^{\text{ap}}}(N)$. For any $k<N$, there are two cases: $A^{\text{ap}}(xy_{:k},x_{k})=0$ and $A^{\text{ap}}(xy_{:k},x_{k})=1$. If $A^{\text{ap}}(xy_{:k},x_{k})=0$ then $y_{k}=1$, and in particular $y_{k}\in h^{\text{am}}(x_{k})$, since $1\in h^{\text{am}}(x)$ for all $x\in X$. If $A^{\text{ap}}(xy_{:k},x_{k})=1$ then $y_{k}=h(x_{k})$, and in particular $y_{k}\in h^{\text{am}}(x_{k})$, since $\{h(x)\}\subseteq h^{\text{am}}(x)$ for all $x\in X$. In either case $y_{k}\in h^{\text{am}}(x_{k})$. Hence, $xy\in\mathcal{C}_{h^{\text{am}}}(N)$.
Consider any $k\in\mathcal{PM}_{h}^{A^{\text{ap}}}(xy)$. Again, there are two cases: $A^{\text{ap}}(xy_{:k},x_{k})=0$ and $A^{\text{ap}}(xy_{:k},x_{k})=1$.
If $A^{\text{ap}}(xy_{:k},x_{k})=0$ then $h(x_{k})=1$ and $0\in A(xy_{:k},x_{k})$. It follows that $h^{\text{am}}(x_{k})=\{1\}$ and $0\in A(xy_{:k},x_{k})\setminus h^{\text{am}}(x_{k})$. In particular, $A(xy_{:k},x_{k})\not\subseteq h^{\text{am}}(x_{k})$ and hence $k\in\mathcal{M}_{h^{\text{am}}}^{A}(xy)$.
If $A^{\text{ap}}(xy_{:k},x_{k})=1$ then $h(x_{k})=0$ and $0\notin A(xy_{:k},x_{k})$. Hence, $y_{k}=0$, because $xy\in\mathcal{C}_{h}^{A^{\text{ap}}}(N)$. In particular, $y_{k}\notin A(xy_{:k},x_{k})$ and hence $k\in\mathcal{M}_{h^{\text{am}}}^{A}(xy)$.
@@
Now, let’s prove Proposition 7.2. The idea is that any frugal ambiguous shattered tree can be interpreted as an adversary strategy for the apple tasting setting. The strategy works by going down the 0 edge when the learner predicts $1$ and going down the $1$ edge (which has to be the $E_{0}$ edge) when the learner predicts 0. In the former case, we always get a mistake (and the edge is always relevant). In the latter case, we get a mistake if and only if the edge is relevant to the ultimate leaf.
Proof of Proposition 7.2.
@@
Construct $xy\in(X\times\{0,1\})^{*}$ and $v\in V^{*}$ recursively as follows. We start from $v_{0}$. Suppose we constructed $xy_{:k}$ and $v_{:(k+1)}$ for some $k\in\mathbb{N}$. If $v_{k}\in V_{f}$, the process ends (i.e. $|xy|=k$ and $|v|=k+1$). Otherwise, set $x_{k}:=\mathbf{x}_{v_{k}}$. Let $z_{k}:=A(xy_{:k},x_{k})$. Set $y_{k}:=1-z_{k}$ and $v_{k+1}:=C_{y_{k}}(v_{k})$.